find the number of years it would take to grow an initial investment of $2,000 to a future value of $5,000, at an interest rate of 8.5%, compounded annually.
formula: FVn=PV(1+i)n
FVn IS THE FUTURE VALUE OF THE INVESTMENT AFTER N YEARS
PV IS THE INITIAL AMOUNT OF THE INVESTMENT
i is the interest rate
n is the length of the investment in years

Found 2 solutions by rfer, mananth:
Answer by rfer(16322)

5000=2000(1.085)^t
5000/2000=1.085^t
2.5=1.085^t
t=11.25 yrs

Answer by mananth(16949)

Amount= 5000
years=n
compounded 1
Rate = 8.5
Amount = P*((n+r)/n)^n

5000 = 2000 *( 1 + 0.09 )^ n
2.5 = *( 1 + 0.09 )^ n
ln 2.5 = n ln 1.09
0.92 = n 0.08
11.23 = n